# Steering AI-Driven Personalization of Scientific Text for General Audiences

### v1

@@ 5.2.2. Participants Who Preferred Lower Degree Liked Concise Translations and Subtle Personalization
Some participants (e.g., P1, P2, P4, P5) preferred translations with lower degrees of personalization primarily because they demanded less mental load to comprehend. They emphasized the cognitive load required to process long personalized translations, stating that visibly longer translations immediately diminished their interest and willingness to read the material (e.g., P2, P5, P13). For instance, translations with lower degrees of personalization tend to be more concise (see Figure 3), which likely contributed to their reduced mental effort.
@@ 5.2.3. Experimenting with Profile Edits
P3: Bar one. I picked the right analogy… And I didn’t mean to. I was just like having fun with it. But that’s a really good analogy, for PFAS I would say.
@@
P6: Oh, I thought it was pretty good. I noticed, specifically, when I changed the profile I liked seeing that drastic differentiation. As you mentioned, age was potentially a factor. I’m glad to see that [age] was taken into consideration in the tone of the message.
@@ 5.3. Impact of Interactive Reading Features on User Experience (RQ2)
We analyze participants’ responses to two features of the tool: generating multiple translations per article and slider interaction to control the personalization degree.
@@ 5.3.1. Compounding Understanding with Multiple Translations
P2: I thought the sugar was BRCA2. It’s actually not. [I noticed from the later translations that] the protein is the BRCA2 and the sugar bypasses the BRCA2.
@@
P3: In real life, I wouldn’t want to read the same thing in a different analogy. If the 1st one didn’t work, I’d be like, ‘Okay, whatever’ you know, I wouldn’t rerun it. I don’t have enough time. [On the other hand] I’ve read different political articles this year, especially with an election year. If there’s something I don’t understand, I’ll go find information elsewhere. In the same way, I would potentially rerun this same interface to see a different analogy or different versions [if I’m interested in the topic].
@@
While participants agreed that the slider was simple and engaging, their desired level of control differed. Some participants expressed interest in having more granular control over the translations (e.g., P3, P7, P8, P9, P10), suggesting that the slider alone might not be sufficient. For example, they wanted to type specific numbers directly to modify the degrees of personalization in addition to slider interaction (e.g., P9) or to be more interactive like ChatGPT (e.g.,  (Laban et al., 2024)).
P10: It would be nice if there was an area of more personalization where I tell them what I like about it, and then they do that [personalization] more or do it less. Like, ‘I would like you [the AI model] to specify less, or I would like you to specify more on either analogy or on the science of it.’
In particular, P10 proposed a feature to allow users to highlight specific parts of the personalized translations and modify them instead of changing the entire translation. They envisioned that this feature could enable users to request more detailed elaboration or varying degrees of personalization for each selected part, providing more granular control over how the tool interprets and modifies different text parts.
In contrast, others preferred minimal control or even automated translations without user intervention at all. Several participants (e.g., P2, P13) found the slider feature overwhelming as it provides too many options (i.e., 100 different versions in the range 0 to 100). For example, P13 suggested 10 or even fewer options for personalization rather than a scale of 0-100. Some proposed having just two versions of the article: original and translated, with an on/off toggle.
Furthermore, P2 suggested a completely automated personalization system. He denied the need for any interactive control options. Rather than investing time or effort in steering various personalization degrees, he envisioned a future where AI would automatically deliver optimally personalized translations based on user preferences and article topics. He believed that AI systems should be capable of determining the optimal degree of personalization for each user and context without manual input. Interestingly, he did not worry about his privacy for such a system as he perceived clear benefits.
@@ 5.4. Potential Use Cases and Concerns
Participants identified the potential use cases and some concerns about this type of tool. Below we describe two primary use cases: learning unfamiliar content and tailoring their own ideas to different audiences. We then present the participants’ two main concerns: reliability of malleable translations and reduced access to the original source.
5.4.1. Learning Unfamiliar Content for Various Motivations (Use Case I)
Participants envisioned using this type of tool to learn unfamiliar topics that they were interested in (e.g., technology (P1), political pledges (P3), economic content (P10), how-to documents (P15)). Particularly, P10 expected that this tool could expand the boundaries of domains that are possible to learn about, as any unfamiliar topic could be explained in a relatable context.
P10: I tend to dodge articles about Econ or business. That’s not my strong suit, so I feel like I would be much more willing to read those. Because usually they’re filled with so much jargon, so like it now opens up a field for me of something that I’m interested in learning.
Furthermore, P3 and P9 envisioned using this tool when they would need to learn something about an unfamiliar topic (e.g., airline compensation policy documents (P3)), which they are not intrinsically motivated to learn, but need to know about (extrinsic motivation).
@@ 5.4.1. Learning Unfamiliar Content for Various Motivations (Use Case I)
Overall, participants found this tool useful to learn about various unfamiliar topics that they wished to explore. Some of them noted that it could help them expand their knowledge into new domains. Moreover, even for topics they found less interesting, the tool could help them quickly grasp essential concepts and key takeaways.
5.4.2. Tailoring Ideas to Communicate to Different Audiences (Use Case II)
One participant (P1) emphasized that this tool could be useful not only to understand unfamiliar topics but also to learn how to communicate complex ideas to diverse audiences. For example, the different degrees of personalization could provide tailored translations of how to explain technical products or business ideas to general audiences. She envisioned saving multiple translations to her personal library and using them during business meetings.
@@ 5.4.2. Tailoring Ideas to Communicate to Different Audiences (Use Case II)
5.4.3. Reliability of Malleable Translations from AI (Concern I)
While most participants focused more on the benefits, one participant (P8) raised concerns about the reliability of AI-generated personalized translations. He argued that personalized translations can vary with each interaction, raising questions about their consistency and trustworthiness.
P8: If I read that I need to use the knowledge, it’s a little more concerning, like I might not trust it as much because it keeps changing. If I had to reference it, or something, I probably wouldn’t trust this. It [AI] might misunderstand something.
He further stated that this concern would motivate him to check out the original paper for complete and accurate understanding. While he would not fully rely on the translations in his professional context, he acknowledged that the tool would be useful in deciding whether to read the original document.
5.4.4. Reduced Access to Original Source Material (Concern II)
One participant (P5) raised a contrasting concern: she was worried that AI-personalized translations might discourage people from accessing original documents. People might rely on these personalized translations rather than engaging with the original material because personalized translations are easier to digest. This could lead to a dependence on AI-processed content at the expense of direct engagement with source materials.
@@ 5.4.4. Reduced Access to Original Source Material (Concern II)
P5 summarized the trade-off associated with this tool. On one hand, it might lead to a larger number of people gaining superficial knowledge across a wider range of topics, without developing a deep understanding or motivation to explore further. On the other hand, without such a tool, it would be less accessible and fewer people would be exposed to or aware of this knowledge in the first place. While acknowledging the concern about shallow understanding, P5 also emphasized increased accessibility to more scientific information.
@@ 6. Discussion
Participants preferred varying degrees of personalized translations. Some preferred concise and less contextual translations (lower degree), while others preferred relatable and contextual ones (higher degree). Interactive reading features enabled them to compound their understanding and even correct initial misunderstandings. Participants noted that these benefits would be particularly valuable when reading topics they were interested in. On the other hand, some participants seeking quick takeaways found little value in exploring multiple translations using the slider. For them, the effort required to explore various translations outweighed the perceived benefits. Given these findings, we discuss several implications of our study in supporting communication in collaborative contexts, maintaining user agency for AI alignment, and the reliability of AI outputs.
6.1. Enhancing Communication by Surfacing Diverse Perspectives
Our findings showed that AI-personalized translations can help users better comprehend scientific content. However, participants perceived broader benefits. They found that personalized translations helped them grasp complex ideas from varied perspectives. The ability to explore multiple viewpoints can prove useful in contexts such as education and cross-domain collaboration.
@@ 6.1. Enhancing Communication by Surfacing Diverse Perspectives
Education has grown increasingly complex, particularly with the widespread adoption of online learning (Li et al., 2021). While online platforms provide greater scalability, instructors face persistent challenges in delivering effective instruction to large, diverse student cohorts (Kulkarni et al., 2015). AI-powered personalization tools can help address these challenges by offering instructors tailored translation suggestions adapted to different students. With appropriate privacy safeguards, these systems could utilize student profiles to recommend more effective concept explanations, enabling more relevant, contextualized interactions between students and instructors. Future research should explore the impact of these personalized translations on learning outcomes.
In cross-disciplinary contexts, effective communication presents similar challenges (Agapie et al., 2022). While cross-domain collaboration is invaluable, several obstacles impede clear communication, such as divergent terminology (Hou and Wang, 2017), varying data interpretations (Kim et al., 2022), and different motivations (Mao et al., 2019). Research shows that new ideas are better understood when presented through familiar examples (Kim et al., 2022). AI-personalized translation tools can bridge these gaps by adapting explanations to collaborators’ backgrounds and expertise, making complex concepts more accessible. This capability is particularly valuable in practical scenarios where cross-disciplinary collaboration is essential for solving complex problems (e.g.,  (Saha et al., 2021; Yoo et al., 2024)). Users can input their collaborators’ domain contexts into the system to receive guidance on effectively communicating their ideas in a way that resonates with their collaborators’ expertise and perspectives.
6.2. Maintaining User Agency in the Pursuit of AI Alignment
Traditionally, the machine learning community has focused on aligning AI models with human intentions through techniques like fine-tuning or reinforcement learning (Ouyang et al., 2022). While these approaches have been effective, they diminish the plurality of cultures and values represented among AI users (Agarwal et al., 2024; Anderson et al., 2024). In this paper, we present an alternative approach that emphasizes user agency. Rather than optimizing the model directly, we designed an interactive tool that empowers users to actively steer and align outputs according to their individual preferences.
This design intervention respects the diversity of experiences and backgrounds of intended general audiences, essential for good science communication (Nisbet and Scheufele, 2009). Our interface provides three key features to support user agency: a) a slider that spans a spectrum of personalization degrees, b) a profile editing function, and c) a history box. These features allow users to influence model outcomes without technical expertise, save personalized preferences, and iteratively explore various outputs until they find desirable results. This approach builds on extensive work in the HCI community which has demonstrated the value of intention-centric design across various domains (e.g.,  (Horvitz, 1999; Amershi et al., 2019)).
@@ 6.2. Maintaining User Agency in the Pursuit of AI Alignment
While highly accurate AI models might better serve diverse user needs, achieving such accuracy often requires extensive user data collection, raising concerns of cost, privacy, and selection bias amongst annotators (Sambasivan et al., 2021). Our design approach provides an alternative: rather than relying on AI-powered systems to perfectly model the plurality of user preferences, we empower users to actively align AI outputs to their individual preferences. This approach not only preserves user privacy but also maintains human agency in human-AI interaction, addressing critical concerns about AI overreliance in our society (Buçinca et al., 2021a; Okolo et al., 2024).
6.3. The Reliability of AI-Personalized Translations
@@
Participants expressed contrasting concerns about the reliability of AI-generated translations for science communication. While some participants were reluctant to rely on the translations, others were worried that people might depend too heavily on them. As AI technology becomes more sophisticated, the higher accuracy of translations with fewer instances of incorrect or hallucinated content might mitigate such reliability concerns (Wu et al., 2017). However, regardless of translation quality, concerns about AI agencies and the trustworthiness of AI may persist. The fact that AI-generated translations are not human-authored continues to affect the perceived trust (Ryan, 2020). This might be due to the unclear responsibility; it is vague who bears the responsibility for mistranslations and their potential consequences. For instance, a participant mentioned that they would not trust AI translations for work-related purposes without verifying the original source, as they would be personally accountable for any errors in the work.
@@
To mitigate risks, we may consider encouraging readers to verify translations against original source materials. However, this might be bothersome. As a participant noted, many people would overly rely on the AI translations and blindly trust what is generated simply because it is convenient, and as we noticed some people do not want to invest mental effort (Buçinca et al., 2021b). This balance between utilizing AI translations and encouraging verification of original sources presents a unique challenge, particularly given the current state of scientific communication. While AI can facilitate making scientific content more accessible, the verification process might be still infeasible as many original scientific articles use overly complex language and jargon, creating barriers for general audiences. Alternatively, we can consider authors (i.e., scientists) to verify several versions of AI-translations before publishing to the general public (Kim et al., 2024). For example, scientists can go through 5-10 different translations with varying degrees of personalization and approve a few of them as verified translations. This would remove the onus of verification from the reader while allowing them to receive the benefits of multiple translations. Further explorations are needed to find scalable approaches such that the verification by authors does not become cumbersome.
6.4. Limitations and Future Studies

### v2

@@ 6.2.2. Participants Who Preferred Lower Degree Liked Concise Translations and Subtle Personalization
Some participants (e.g., P1, P2, P4, P5) preferred translations with lower degrees of personalization primarily because they demanded less mental load to comprehend. They emphasized the cognitive load required to process long personalized translations, stating that visibly longer translations immediately diminished their interest and willingness to read the material (e.g., P2, P5, P13). For instance, translations with lower degrees of personalization tend to be more concise (Figure 5), which likely contributed to their reduced mental effort.
@@ 6.2.3. Experimenting with Profile Edits
P3: Bar one. I picked the right analogy… And I didn’t mean to. I was just like having fun with it. But that’s a really good analogy, for PFAS [Per- and Polyfluoroalkyl Substances] I would say.
@@
P6: Oh, I thought it was pretty good. I noticed, specifically, when I changed the profile I liked seeing that drastic differentiation. As you mentioned, age was potentially a factor. I’m glad to see that [age] was taken into consideration in the message tone.
@@ 6.3. Impact of Interactive Reading Features on User Experience (RQ2)
We analyze participants’ responses to two tool features: generating multiple translations per article and slider interaction to control the personalization degree.
@@ 6.3.1. Compounding Understanding with Multiple Translations
P2: I thought the sugar was BRCA2 [BReast CAncer gene 2]. It’s actually not. [I noticed from the later translations that] the protein is the BRCA2 and the sugar bypasses the BRCA2.
@@
P3: In real life, I wouldn’t want to read the same thing in a different analogy. If the 1st one didn’t work, I’d be like, ‘Okay, whatever’ you know, I wouldn’t rerun it. I don’t have enough time. However, I’ve read different political articles this year, especially with an election year. If there’s something I don’t understand, I’ll go find information elsewhere. In the same way, I would potentially rerun this to see a different analogy or different versions [if I’m interested in the topic].
@@
While participants agreed that the slider was simple and engaging, their desired level of control differed. Some participants expressed interest in having more granular control over the translations (e.g., P3, P7, P8, P9, P10), suggesting that the slider alone might not be sufficient. For example, they wanted to manually type specific numbers to modify the degrees of personalization in addition to slider interaction (e.g., P9) or prompt interactions to specify more or less personalization.
P10: It would be nice if there were an area of more personalization where I tell them what I like about it, and then they do that [personalization] more or do it less. Like, ‘I would like you [the AI model] to specify less, or I would like you to specify more on either analogy or on the science of it.’
In particular, P10 proposed a feature to allow users to highlight specific parts of the personalized translations and modify them instead of changing the entire translation (e.g., selective editing features (Laban et al., 2024)). They envisioned that this feature could enable users to request more detailed elaboration or varying degrees of personalization for each selected part, providing more granular control over how the tool interprets and modifies different text parts.
In contrast, others preferred minimal control or even automated translations without any user intervention. Several participants (e.g., P2, P13) found the slider feature overwhelming as it provides too many options (i.e., 100 different versions in the range 0 to 100). For example, P13 suggested 10 or even fewer options for personalization rather than a scale of 0-100. Some proposed having just two versions of the article: original and translated, with an on/off toggle.
Furthermore, some participants (e.g., P2) suggested a completely automated personalization system. Rather than steering various personalization degrees, they wanted the AI to automatically deliver optimally personalized translations based on user preferences and article topics. Further, the perceived benefits from the system outweighted privacy concerns.
@@ 6.4. Potential Use Cases and Concerns
Participants identified several potential use cases and some concerns. Below, we highlight two use cases: learning unfamiliar content and tailoring their ideas to different audiences. We then present two concerns: the reliability of malleable translations and reduced access to the source article.
6.4.1. Use Case I: Learning Unfamiliar Content for Various User Motivations
Participants envisioned using this type of tool to learn unfamiliar topics that they were interested in [intrinsic motivation] such as technology (P1), political pledges (P3), economic content (P10), and how-to documents (P15). Particularly, P10 envisioned using this tool for expanding their knowledge boundaries by learning about unfamiliar or new domains.
P10: I tend to dodge articles about Econ and business. They are not my strong suit. I feel like I would be much more willing to read those [through TranSlider]. Because usually they’re filled with so much jargon, so like it now opens up a field for me of something that I’m interested in learning.
Furthermore, P3 and P9 envisioned using this tool to learn something about an unfamiliar topic (e.g., airline compensation policy documents (P3)), which they are not intrinsically motivated to learn, but need to know about [extrinsic motivation].
@@ 6.4.1. Use Case I: Learning Unfamiliar Content for Various User Motivations
Overall, participants viewed TranSlider as a learning tool to explore various unfamiliar topics and new domains. Moreover, for topics they found less interesting, the tool could help them quickly grasp essential concepts and key takeaways.
6.4.2. Use Case II: Tailoring Ideas to Communicate to Different Audiences
Participants (e.g., P1) also envisioned using this tool to learn how to communicate complex ideas to diverse audiences. For example, the different degrees of personalization could provide tailored translations of how to explain technical products or business ideas to general audiences. P1 envisioned saving multiple translations to their personal library and using them during business meetings.
@@ 6.4.2. Use Case II: Tailoring Ideas to Communicate to Different Audiences
6.4.3. Concern I: Reliability of Malleable Translations from AI
While most participants focused more on the benefits, one participant (P8) raised concerns about the reliability of AI-generated personalized translations. They argued that personalized translations can vary with each interaction, raising questions about their consistency and trustworthiness.
P8: If I read something and then have to apply that information in my work, it’s a little more concerning. I might not trust it because it keeps changing. If I had to cite it, I probably wouldn’t trust this. It [AI] might misunderstand something.
They further stated that this concern would motivate them to check out the original paper for a complete and accurate understanding. While they would not fully rely on the translations in their professional context, they acknowledged that the tool would be useful in deciding whether to read the original document.
6.4.4. Concern II: Reduced Reference to Original Source Material
One participant (P5) raised a contrasting concern: she worried that AI-personalized translations might discourage people from referring to original documents. People might rely on these personalized translations rather than engaging with the original material because personalized translations are easier to digest. This could lead to a dependence on AI-processed content at the expense of direct engagement with source materials.
@@ 6.4.4. Concern II: Reduced Reference to Original Source Material
P5 summarized the trade-off: the tool could broaden superficial knowledge access across wider ranging topics but hinder deep understanding and further exploration. Conversely, P5 weighed the risk of shallow learning against the benefit of increased accessibility, noting that without the tool, fewer individuals would gain any exposure to such knowledge.
@@ 7. Discussion
7.1. Implications for Science Communication and the CSCW Community
Our study demonstrated the feasibility of addressing the translation challenge in science communication. Rather than pursuing a single ‘optimal’ translation, TranSlider generates and enables exploring multiple personalized translations. This approach produced two key benefits: i) a compounding understanding of scientific content, where each translation contributed like a puzzle piece to build a more complete picture of the underlying content, and ii) enabling multi-perspective comprehension, where varied viewpoints conveyed different aspects of the content.
7.1.1. Towards Collective Understanding of Science with General Audiences
Our approach could address the dissemination issue in science communication: scientists’ reluctance to disseminate their work on digital media. Their concerns stem from the potential misinformation arising during translation, which could make audiences misinterpret simplified explanations (Wu et al., 2019; Loroño-Leturiondo and Davies, 2018). By enabling readers to explore multiple personalized translations, the TranSlider might create an opportunity for them to build a comprehensive understanding while potentially reducing the risk of misinformation. This approach aligns with established science communication practices, where journalists often craft multiple re-expressions when content might be unfamiliar to general audiences (Hullman et al., 2018).
We envision collective knowledge sharing spaces inspired by existing AI-generated content platforms, such as Midjourney (Mid, 2025) and DeviantArt (dev, 2025). The Midjourney system shares users’ generated images throughout its community by default, allowing individuals to inspire each other to create new and unique creative works along with constructive discussions within the community (Guo et al., 2024). Applying similar dynamics to science communication, platforms like Google Scholar (Goo, 2025) or Semantic Scholar (Sem, 2025) could integrate our approach to let general audiences explore scientific articles and share personalized translations across their community spaces (e.g.,  OtherTube (Bhuiyan et al., 2022)). Creating such environments could foster diverse perspectives and collaborative comprehension of complex scientific materials, and further the larger vision of making science accessible at scale.
Such collective community spaces could also create valuable opportunities for scientists to observe how general audiences digest and interact with their research. Traditionally, scientists have limited visibility into the public perception and potential implications of their work, largely due to a lack of active engagement with broader audiences (Williams et al., 2022). However, within this community setting, researchers could preview how individual and collective understandings develop around their work. These insights could influence or engender new research directions. Additionally, such feedback might help scientists identify and even prevent potentially problematic research activities that they would not otherwise anticipate without such collective reflection through public interpretation (Llorente et al., 2019).
However, designers should be aware that malicious actors could exploit such community ecosystems to manipulate the collective understanding of certain topics. Furthermore, some organizations (e.g.,  ‘Big Brother’ (Orwell, 2024)) might leverage these platforms to influence science activities, as public interest often drives funding allocation, and without proper financial support, some research endeavors may not remain sustainable. We see this line of research as a promising new direction.
7.1.2. Enhancing Communication by Surfacing Diverse Perspectives
@@
The benefit of having varied perspectives could be useful in contexts such as education and cross-disciplinary collaboration. Education has grown increasingly complex, particularly with the widespread adoption of online learning (Li et al., 2021). While online platforms provide greater scalability, instructors face persistent challenges in delivering effective instruction to large, diverse student cohorts (Kulkarni et al., 2015). AI-powered personalization tools could help address these challenges by offering instructors tailored translation suggestions adapted to different students. For example, systems could let children input their favorite animals, foods, or objects relevant to the given learning context, and generate tailored explanations and learning materials to better engage students. Future research should explore the impact of these personalized translations on learning outcomes, such as learning performance.
Likewise, in cross-disciplinary contexts, effective communication is often challenging due to several obstacles, such as divergent terminology (Hou and Wang, 2017), varying data interpretations (Kim et al., 2022), and different motivations (Mao et al., 2019). Research shows that new ideas are better understood when presented through familiar examples (Kim et al., 2022). Cross-disciplinary collaborative activities (e.g., drafting documents, building systems, and shaping policy letters) require clear and constant communication among collaborators (Saha et al., 2021; Yoo et al., 2024). While real-time conversations offer opportunities to address misunderstandings immediately, scheduling synchronous meetings among collaborators inevitably creates communication delays that can slow project progress. Our interface could benefit such asynchronous communication scenarios by allowing collaborators to generate personalized translations of complex content, helping team members from different disciplinary backgrounds understand shared documents without waiting for clarification meetings.
However, these approaches may lead to over-reliance on AI-personalization, as indicated by a participant (P5). In a future where AI tools work exceptionally well at addressing communication barriers, individuals might prefer interacting with these interfaces rather than directly communicating with students, instructors, or collaborators. The convenience of AI-generated ‘perfect’ translations might even diminish individuals’ efforts and motivation to develop clear communication skills and articulate their ideas in accessible ways. This raises crucial concerns about balancing between technological assistance and the development of fundamental human agency in communication capabilities, suggesting that such tools should be designed to augment rather than replace human roles. Further research is warranted on how to design and develop such AI-driven cross-disciplinary collaborative interfaces while balancing the perceived benefits, potential long-term harms, and ethical concerns.
7.2. Interface Design Enabling User Control for Human-AI Alignment
We designed an interactive interface that empowers users to steer and align AI outputs to their preferences, as opposed to optimizing the AI models directly (Ouyang et al., 2022). While many participants appreciated the controllability of AI output, we found some users may not want such control in-the-wild. Participants also showed varied responses about the reliability of AI outputs.
7.2.1. Flexible AI Interfaces for Balancing User Control and Automation
@@
Our interface provides three key features to support user agency: a) a slider that allows users to adjust the degree of personalization, b) a profile editing function, and c) a history box. These features enable users to control AI outputs without technical expertise, save personalized preferences, and iteratively explore various outputs until they find desirable results. While highly accurate AI models might better serve diverse user needs, achieving such accuracy often requires extensive user data collection, raising concerns of cost, privacy, and selection bias amongst annotators (Sambasivan et al., 2021). Our design approach provides an alternative: rather than relying on AI-powered systems to perfectly model the plurality of user preferences, we empower users to actively align AI outputs to their individual preferences. This approach not only preserves user privacy but also maintains human agency in human-AI interaction, addressing critical concerns about AI overreliance in our society (Buçinca et al., 2021; Okolo et al., 2024).
However, we should acknowledge a potential limitation of our approach. Our interaction design assumes users want to preserve their agency over AI systems, which may not always be true depending on individuals and contexts. Indeed, a participant (P2) stated they do not want to interact with systems in-the-wild (Section 6.3.2). As our study suggests, in designing systems for mass audiences, we must recognize that a portion of users, like P2, may prefer fully automatic personalization without manual adjustments. Also, in different contexts such as entertainment and creative art domains, users may prefer consuming the contents passively, rather than proactively interacting with systems (Kim et al., 2024). Therefore, future systems should balance automation and user control, perhaps by offering flexible interfaces (e.g.,  (Min et al., 2025)) providing both automatic and control modes that users can switch between depending on their needs.
7.2.2. The Reliability of AI-Personalized Translations
@@
Participants expressed contrasting concerns about the reliability of AI-generated translations for science communication. While some participants were reluctant to rely on the translations, others were worried that people might depend too heavily on them. As AI technology becomes more sophisticated, the higher accuracy of translations with fewer instances of incorrect or hallucinated content might mitigate such reliability concerns (Wu et al., 2017). However, regardless of translation quality, concerns about AI-processed output and the trustworthiness of AI may persist. The fact that AI-generated translations are not human-authored continues to affect the perceived trust (Ryan, 2020). The perceived mistrust could stem from an ambiguity in responsibility, specifically regarding the onus for mistranslations and their resultant ramifications. For instance, a participant mentioned that they would not trust AI outputs for work-related tasks without verifying the source, as they are personally accountable for any errors in the work.
@@
To mitigate risks, verifying translations against original source materials could be an approach. However, who should bear the responsibility: readers or authors? Putting the onus on the reader defeats the purpose of personalization for making science accessible as original scientific articles use overly complex language and jargon. Our findings and prior work indicate that readers do not want to invest significant mental effort parsing complex texts, resulting in the risk of over reliance on AI-generated translations (Buçinca et al., 2021). A potential solution could be for authors (i.e.,  scientists) to verify several versions of AI translations before publishing to the general public (Kim et al., 2024). For example, scientists could go through 5–10 different translations with varying degrees of personalization and approve a subset as verified translations. The balance between utilizing AI translations and encouraging verification of original sources presents a unique challenge, particularly given the current state of scientific communication. Further explorations are needed to find scalable approaches such that the author verification does not become cumbersome.
7.3. Limitations and Future Work